You have a bag of marbles. There are twice as many blue marbles as red marbles. There are 4 less green marbles than red marbles. If there are 28 marbles in the bag, how many of each marble is in the bag. Use the 5 step solution, please.
Answer by stanbon(75887)

You have a bag of marbles. There are twice as many blue marbles as red marbles. There are 4 less green marbles than red marbles. If there are 28 marbles in the bag, how many of each marble is in the bag. Use the 5 step solution, please.
Equations:
b + r + g = 28
b = 2r
g = r-4
-----
Substitute and solve for "r":
2r + r + r-4 = 28
4r = 32
r = 8 red
b = 2r = 16 blue
g = r-4 = 4 green
====================
Cheers,
Stan H.